# Masked Generative Extractor for Synergistic Representation and 3D Generation of Point Clouds

### v1

@@ 3.4 Iterative Generation
After pretraining, following Li_2023_CVPR , we employ an iterative generation strategy to generate diverse 3D shapes. The generation process starts with all tokens being considered as masked tokens. Using a cosine function, we determine the number of masked tokens to be replaced in each iteration. Specifically, in each iteration, based on the predicted probabilities from the generator, we sample and replace those tokens with higher probabilities.
Table 1: Shape classification results on ModelNet40 Wu_2015_CVPR and ScanObjectNN Uy_2019_ICCV . We report the accuracy (%) of both training from scratch and fine-tuning models on ModelNet40 and ScanObjectNN datasets.
Methods Reference ModelNet40 ScanObjectNN
OBJ_BG OBJ_ONLY PB_T50_RS
training from scratch
PointNet Qi_2017_CVPR CVPR 2017 89.2 73.3 79.2 68.0
PointNet++ qi2017pointnet++ Neurips 2017 90.5 82.3 84.3 77.9
PointCNN li2018pointcnn Neurips 2018 92.2 86.1 85.5 78.5
DGCNN wang2019dynamic TOG 2019 92.9 82.8 86.2 78.1
GBNet 9410405 TMM 2021 93.8 - - -
MVTN Hamdi_2021_ICCV ICCV 2021 93.8 92.6 - 82.8
PointNeXt qian2022pointnext Neurips 2022 94.0 - - -
fine-tuning
Point-BERT Yu_2022_CVPR CVPR 2022 93.2
Mask-Point liu2022masked ECCV 2022 93.8 89.3 88.1 84.3
Point-MAE pang2022masked ECCV 2022 93.8 90.0 88.2 85.2
Point-M2AE zhang2022point Neurips 2022 94.0 91.2 88.8 86.4
PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 94.0 91.6 90.0 86.9
Point-MAGE - 94.2 92.9 91.0 87.1
Table 2: Few-shot classification on ModelNet40Wu_2015_CVPR . We report the average accuracy (%) and standard deviation (%) of 10 independent experiments.
Methods Reference 5-way 10-way
@@
training from scratch
DGCNN wang2019dynamic TOG 2019 31.6 ± 2.8 40.8 ± 4.6 19.9 ± 2.1 16.9 ± 1.5
OcCo Wang_2021_ICCV ICCV 2021 90.6 ± 2.8 92.5 ± 1.9 82.9 ± 1.3 86.5 ± 2.2
fine-tuning
Transformer + OcCo Yu_2022_CVPR CVPR 2022 94.0 ± 3.6 95.9 ± 2.3 89.4 ± 5.1 92.4 ± 4.6
Point-BERT Yu_2022_CVPR CVPR 2022 94.6 ± 3.1 96.3 ± 2.7 91.0 ± 5.4 92.7 ± 5.1
Mask-Point liu2022masked ECCV 2022 95.0 ± 3.7 97.2 ± 1.7 91.4 ± 4.0 93.4 ± 3.5
Point-MAE pang2022masked ECCV 2022 96.3 ± 2.5 97.8 ± 1.8 92.6 ± 4.1 95.0 ± 3.0
Point-M2AE zhang2022point Neurips 2022 96.8 ± 1.8 98.3 ± 1.4 92.3 ± 4.5 95.0 ± 3.0
PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 96.8 ± 2.0 98.6 ± 1.1 92.6 ± 4.6 95.2 ± 3.4
Point-MAGE - 97.2 ± 1.9 98.6 ± 0.7 92.7 ± 4.0 95.4 ± 3.0
4 Experiments
In this section, we first provide a detailed overview of the pretraining setup of Point-MAGE. We then evaluate the pretrained model’s performance in both representation learning and generation capabilities. Finally, to gain a deeper understanding of the roles and importance of each component in the model, we conduct ablation studies on Point-MAGE. More results can be found in the Appendix.
@@ 4.1 Pre-training Setups
Point-MAGE is pre-trained on the ShapeNet chang2015shapenet dataset, which includes 55 different object categories and over 50,000 3D models. From the surface of each 3D shape, 2048 points are uniformly sampled and further divided into 512 patches, each containing 32 points. In the first stage, we pre-trained the VQVAE to obtains a discrete token for each point patch, resulting in a token sequence of 512 tokens. In the second stage, we determine the variable masking ratio with $\mu=0.55$ and $\sigma=0.25$ and pre-train the extractor-generator architecture based on masked token reconstruction. Additionally, to validate the capability of integrating the Point-MAGE framework with advanced model designs, we pre-trained a multi-scale extractor model following zhang2022point and evaluated its representation learning performance across various downstream tasks. We used the AdamW optimizer for pre-training the two stages for 800 and 300 epochs, respectively. Further implementation details and hyperparameter settings are provided in the Appendix.
@@ 4.2 Representation Learning
To demonstrate the representation learning performance of our method, we conducted experiments on shape classification, few-shot learning, and part segmentation. Additionally, the Appendix will provide further research and discussion on the properties of Point-MAGE.
@@
We conducted shape classification fine-tuning on two datasets. The ModelNet40 Wu_2015_CVPR dataset includes 3D CAD models of 40 categories, and we followed the official data split provided in  Wu_2015_CVPR for our experiments. As shown in Table 1, Point-MAGE achieved a classification accuracy of 94.2% on this dataset, surpassing other existing pre-trained models. The ScanObjectNN  Uy_2019_ICCV dataset, derived from real-world indoor scans, contains approximately 15,000 objects across 15 categories. We evaluated this dataset using three different data splits: OBJ-BG, OBJ-ONLY, and PB-T50-RS. The experimental results indicate that Point-MAGE consistently outperformed other models across all splits. Notably, in the OBJ-BG setting, Point-MAGE achieved an accuracy of 92.9%, outperforming the second-best pre-trained model NEURIPS2023_5ed5c3c8 by 1.3 percentage points.
@@
We evaluated the few-shot learning performance of Point-MAGE by conducting N-way, K-shot experiments on the ModelNet40Wu_2015_CVPR dataset. Specifically, we randomly selected N classes from ModelNet40 and sampled K objects from each class. Each experiment was repeated 10 times, and the average and standard deviation of the results were calculated. We evaluated each of the four combinations of settings with N = {5, 10} and K = {10, 20}. As shown in the table 2, we achieved SOTA performance in all of them, demonstrating that Point-MAGE’s pre-trained 3D representation is well-suited for few-shot learning tasks.
Table 3: Part segmentation on ShapeNetPart yi2016scalable . We report the mean IoU (%) of all part categories (’$\text{mIoU}_{C}$’) and the mean IoU of all instances (’$\text{mIoU}_{I}$’) in the dataset.
       Methods        Reference        $\text{mIoU}_{C}$uparrow        $\text{mIoU}_{I}$uparrow
       training from scratch
       PointNet Qi_2017_CVPR        CVPR 2017        80.4        83.7
       PointNet++ qi2017pointnet++        Neurips 2017        81.9        85.1
       DGCNN wang2019dynamic        TOG 2019        82.3        85.2
       fine-tuning
       Transformer + OcCo Yu_2022_CVPR        CVPR 2022        83.4        85.1
       Point-BERT Yu_2022_CVPR        CVPR 2022        84.1        85.6
       Mask-Point liu2022masked        ECCV 2022        84.4        86.0
       Point-MAE pang2022masked        ECCV 2022        -        86.1
       Point-M2AE zhang2022point        Neurips 2022        84.9        86.5
       PointGPT NEURIPS2023_5ed5c3c8        Neurips 2023        84.1        86.2
       Point-MAGE        -        85.0        86.5
@@
Methods Surface-FPD $\downarrow$ Surface-KPD (×$10^{3}$$\downarrow$ Rendering-FID  $\downarrow$ Rendering-KID (×$10^{3}$$\downarrow$
Grid-$8^{3}$ zhang20223dilg 4.03 6.15 32.78 14.12
3DILG zhang20223dilg 1.89 2.17 24.83 10.51
Point-MAGE 0.98 1.34 22.79 11.48
Figure 2: Visualization of unconditional generation. All models were trained on the complete ShapeNet dataset, and we sampled nine generated results for rendering for each model.
Part Segmentation.
Table 3 presents the results of average IoU for all categories ($\text{mIoU}{C}$) and all instances ($\text{mIoU}{I}$) in the part segmentation experiments on the ShapeNetPart yi2016scalable dataset. Compared to the fully supervised method DGCNNwang2019dynamic , Point-MAGE achieved significant improvements of 2.7% and 1.3% in these two key evaluation metrics, respectively. Our method also outperforms other pretrained models, achieving a $\text{mIoU}_{C}$ of 85.0%, which ranks first.
4.3 3D Shape Generation
@@
After completing the pre-training stage, Point-MAGE is capable of performing unconditional generation tasks. Fréchet Inception Distance (FID) and Kernel Inception Distance (KID) are commonly used to evaluate the generation quality of generative models. Following zhang20233dshape2vecset , we use Rendering-FID and Rendering-KID metrics to evaluate the FID and KID of rendered images of 3D shapes, as well as Surface-FPD and Surface-KPD metrics to evaluate the FID and KID of sampled point clouds from 3D shapes. Table 4 compares the performance of various methods in unconditional shape generation on ShapeNet. 3DILG also adopts a two-stage generation strategy and completes the generation training through an autoregressive process. Point-MAGE outperforms other methods on all three evaluation metrics, indicating that our method is capable of generating 3D shapes that match the real distribution. Additionally, we visualize the rendered results of generated shapes for all methods in Figure 2.
@@
We evaluated the conditional generation capabilities of Point-MAGE under two different settings. First, we retrained the model using the 54 categories from the ShapeNet dataset as conditional embeddings. Figure 3 displays the rendered results of generated shapes for three categories (airplane, faucet, and table), with nine instances shown for each category. These visualizations demonstrate our method’s ability to generate realistic 3D shapes based on categorical conditions. Next, we further evaluated Point-MAGE against other SSL models in the masked generation task. As shown in Figure 4, these models typically train by reconstructing local point clouds, which primarily allows them to complete point cloud reconstruction tasks and depends on the provision of the central point positions. This approach, to some extent, leaks complete shape information to the decoder. In contrast, our method can reconstruct 3D shape meshes using the minimal amount of point cloud information and does not rely on masked center point information during the generation process.
@@ 4.4 Ablation Studies
In this section, we discuss three aspects that influence the representation learning and generation capabilities of the pre-trained model: the transformer design of the extractor-generator architecture, the impact of different inputs, and the effect of varying masking ratios.
@@
As shown in Table 5(a), we used the standard ViT architecture to build the extractor-generator structure and explored four different transformer depth combinations. The results indicate that increasing the model depth significantly enhances the performance of the pre-trained model. Additionally, consistent with trends observed in previous masked modeling studies he2022masked ; pang2022masked ; zhang2022point , an asymmetric extractor-generator structure is more beneficial for representation learning. This reliance on the extractor to obtain high-quality representations from a small number of visible tokens is due to the use of a relatively simple generator to complete the reconstruction task.
@@
In Table 5(b), we investigate the impact of two different input types on model performance. Specifically, after vector quantizing the initial features of the point cloud patches, we consider using either the quantized features or their indices in the codebook as inputs. Experimental results show that using index-embedded tokens benefits the generation task, but its generalization ability across datasets is limited, which is detrimental to fine-tuning on downstream tasks, leading to poorer representation learning capabilities. This contrasts sharply with the results shown by MAGE Li_2023_CVPR , where using index embeddings performs well in both aspects. We believe this characteristic of MAGE arises from its pre-training on a larger-scale dataset, and the advantages of image data over point cloud data in terms of regularity and consistency, which reduce the impact of different data sources.
@@
To determine the optimal masking ratio, in Table 5(c), we compare the impact of various masking ratio distributions on fine-tuning and near-unconditional generation task results. Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the truncated Gaussian distribution, respectively. When $\sigma=0$, a fixed masking ratio is used. In this case, when $\mu=1.0$, all tokens are masked, preventing the extractor from obtaining useful information, making the fine-tuning process essentially equivalent to training from scratch. In other cases, since the generator does not experience the process of reconstructing from zero, it fails to complete the generation task, and the generation results are marked as not applicable (N/A). The findings indicate that employing a variable masking ratio not only enables the completion of generation tasks but also encourages the representation learning process to capture superior features.
Table 5: Ablation experiments. The fine-tuned accuracy on the ScanObjectNN Uy_2019_ICCV (OBJ_BG) and the Rendering-FID (R-FID) for unconditional generation are reported.
(a) Transformer depth
Extractor Generator Acc (%) R-FID $\downarrow$
6 6 91.1 29.07
8 6 91.7 27.85
12 8 92.4 22.79
12 12 91.9 20.94
(b) Input type
Type Acc (%) R-FID $\downarrow$
Quantized features 92.4 22.79
Index-embedded tokens 87.3 22.36
(c) Masking ratio distribution
$\mu=1.0$ $\mu=0.7$ $\mu=0.6$ $\mu=0.55$ $\mu=0.5$ $\mu=0.55$ $\mu=0.55$ $\mu=0.55$
$\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0.15$ $\sigma=0.25$ $\sigma=0.5$
Acc (%) 86.2 91.8 91.9 92.0 91.8 92.2 92.4 92.1
R-FID $\downarrow$ 26.72 N/A N/A N/A N/A 23.58 22.79 23.65

### v2

@@ 3.4 Iterative Generation
After pretraining, following Li et al. (2023), we employ an iterative generation strategy to generate diverse 3D shapes. The generation process starts with all tokens being considered as masked tokens. Using a cosine function, we determine the number of masked tokens to be replaced in each iteration. Specifically, in each iteration, based on the predicted probabilities from the generator, we sample and replace those tokens with higher probabilities.
4 Experiments
In this section, we first provide a detailed overview of the pretraining setup of Point-MGE. We then evaluate the pretrained model’s performance in both representation learning and generation capabilities. Finally, to gain a deeper understanding of the roles and importance of each component in the model, we conduct ablation studies on Point-MGE. More results can be found in the Appendix.
Methods 5-way 10-way
@@ 4 Experiments
Supervised Learning Only
PointNet 52.0±3.8 57.8±4.9 46.6±4.3 35.2±4.8
PointNet-OcCo 89.7±1.9 92.4±1.6 83.9±1.8 89.7±1.5
DGCNN 31.6±2.8 40.8±4.6 19.9±2.1 16.9±1.5
with Single-Modal Pre-Training
Point-BERT 94.6±3.1 96.3±2.7 91.0±5.4 92.7±5.1
Mask-Point 95.0±3.7 97.2±1.7 91.4±4.0 93.4±3.5
Point-MAE 96.3±2.5 97.8±1.8 92.6±4.1 95.0±3.0
Point-M2AE 96.8±1.8 98.3±1.4 92.3±4.5 95.0±3.0
PointGPT 96.8±2.0 98.6±1.1 92.6±4.6 95.2±3.4
GPM 97.2±2.6 98.7±2.2 92.9±4.2 95.0±3.0
Point-MGE 97.2±1.7 98.7±0.7 92.8±4.0 95.4±3.2
Improvement +2.6 +2.4 +1.8 +2.7
PointGPT-B 97.5±2.0 98.8±1.0 93.5±4.0 95.8±3.0
Point-MGE-B 97.7±1.8 98.8±0.8 93.8±4.6 95.8±3.4
with Cross-Modal Pre-Training
ACT 96.8±2.3 98.0±1.4 93.3±4.0 95.6±2.8
Joint-MAE 96.7±2.2 97.9±1.8 92.6±3.7 95.1±2.6
I2P-MAE 97.0±1.8 98.3±1.3 92.6±5.0 95.5±3.0
Recon 97.3±1.9 98.9±1.2 93.3±3.9 95.8±3.0
Table 2: Few-shot classification on ModelNet40 Wu et al. (2015). We report the average accuracy (%) and standard deviation (%) of 10 independent experiments.
@@ 4.1 Pre-training Setups
Point-MGE is pre-trained on the ShapeNet Chang et al. (2015) dataset, which includes 55 different object categories and over 50,000 3D models. From the surface of each 3D shape, 2048 points are uniformly sampled and further divided into 64 patches, each containing 32 points. First, we trained the VQVAE to obtains the discrete tokens. Next, we determined the sliding masking ratio, setting $\beta=0.5$ and $u=2$, and pre-trained the ViT-based extractor-generator architecture using multi-scale masked token reconstruction. The number of scales and the number of point patches at each scale were consistent with Point-M2AE Zhang et al. (2022b). Additionally, to validate the high-capacity model learning ability of the Point-MGE framework, we scaled the model to ViT-B as described in Chen et al. (2023), performed post-training, and evaluated its representation learning performance across various downstream tasks. We used the AdamW optimizer to perform pre-training for 300 epochs. Further implementation details and hyperparameter settings are provided in the Appendix.
@@ 4.2 Representation Learning
To demonstrate the representation learning performance of our method, we conducted experiments on shape classification, few-shot learning, and part segmentation. We evaluated the benchmark performance for supervised learning only, single-modal pre-training, and cross-modal pre-training separately.
@@
We conducted shape classification fine-tuning on two datasets. The ModelNet40 Wu et al. (2015) dataset includes 3D CAD models of 40 categories, and we followed the official data split provided in  Wu et al. (2015) for our experiments. As shown in Table 1, Point-MGE achieved a second-place classification accuracy of 94.2% on this dataset. The ScanObjectNN  Uy et al. (2019)dataset, derived from real-world indoor scans, contains approximately 15,000 objects across 15 categories. We evaluated this dataset using three different data splits: OBJ-BG, OBJ-ONLY, and PB-T50-RS. The experimental results indicate that Point-MGE consistently outperformed other models across all splits. Notably, in the OBJ-BG setting, Point-MGE achieved an accuracy of 92.9%, outperforming the second-best pre-trained model Chen et al. (2023) by 1.3 percentage points. Additionally, we extended Point-MGE and three benchmark methods to the ViT-B model for performance evaluation. The results demonstrate that under high-capacity model conditions, our method still maintains a competitive advantage.
Methods $\text{mIoU}_{C}$uparrow $\text{mIoU}_{I}$uparrow
Supervised Learning Only
PointNet Qi et al. (2017a) 80.4 83.7
PointNet++ Qi et al. (2017b) 81.9 85.1
DGCNN Wang et al. (2019) 82.3 85.2
with Single-Modal Pre-Training
Transformer + OcCo Yu et al. (2022) 83.4 85.1
Point-BERT Yu et al. (2022) 84.1 85.6
Mask-Point Liu et al. (2022) 84.4 86.0
Point-MAE Pang et al. (2022) - 86.1
PointGPT Chen et al. (2023) 84.1 86.2
GPM Li et al. (2024) 84.2 85.8
Point-MGE 85.0 86.5
Improvement +0.9 +0.9
PointGPT-B Chen et al. (2023) 84.5 86.5
Point-MGE-B 85.0 86.6
with Cross-Modal Pre-Training
ACT Dong et al. (2023) 84.7 86.1
Recon Qi et al. (2023) 84.8 86.4
Table 3: Part segmentation on ShapeNetPart Yi et al. (2016). We report the mean IoU (%) of all part categories (’$\text{mIoU}_{C}$’) and the mean IoU of all instances (’$\text{mIoU}_{I}$’) in the dataset.
@@
We evaluated the few-shot learning performance of Point-MGE by conducting N-way, K-shot experiments on the ModelNet40Wu et al. (2015) dataset. Specifically, we randomly selected N classes from ModelNet40 and sampled K objects from each class. Each experiment was repeated 10 times, and the average and standard deviation of the results were calculated. We evaluated each of the four combinations of settings with N = {5, 10} and K = {10, 20}. As shown in Table 2, under both ViT and ViT-B settings, we achieved SOTA performance in at least three combinations, demonstrating that Point-MGE’s pre-trained 3D representations are highly effective for few-shot learning tasks.
Part Segmentation.
Table 3 presents the results of average IoU for all categories ($\text{mIoU}_{C}$) and all instances ($\text{mIoU}_{I}$) in the part segmentation experiments on the ShapeNetPart Yi et al. (2016) dataset. Compared to the fully supervised method DGCNNWang et al. (2019), Point-MGE achieved significant improvements of 2.7% and 1.3% in these two key evaluation metrics, respectively. Our method also outperforms other pretrained models, achieving a $\text{mIoU}_{C}$ of 85.0%, which ranks first. Additionally, the experimental results show that increasing the number of parameters yields only marginal improvements for this task.
4.3 3D Shape Generation
Methods S-FPD $\downarrow$ S-KPD $\downarrow$ R-FID $\downarrow$ R-KID $\downarrow$
(×$\times 10^{3}$) (×$\times 10^{3}$)
Grid-$8^{3}$ 4.03 6.15 32.78 14.12
PVD 2.33 2.65 270.64 281.5
3DILG 1.89 2.17 24.83 10.51
3DShape2VecSet 0.97 1.11 24.24 11.76
Point-MGE 0.85 1.34 21.63 11.48
@@
Figure 2: Visualization of unconditional generation. All models were trained on the complete ShapeNet dataset, and we sampled nine generated results for rendering.
@@
After completing the pre-training stage, Point-MGE is capable of performing unconditional generation tasks. fréchet inception distance (FID) and kernel inception distance (KID) are commonly used to evaluate the generation quality of generative models. Following Zhang et al. (2023a), we use Rendering-FID (R-FID) and Rendering-KID (R-KID) metrics to evaluate the FID and KID of rendered images of 3D shapes, as well as Surface-FPD (S-FPD) and Surface-KPD (S-KPD) metrics to evaluate the FID and KID of sampled point clouds from 3D shapes. Table 4 compares the performance of various methods in unconditional shape generation on ShapeNet. 3DILG Zhang et al. (2022a) also adopts a two-stage generation strategy and completes the generation training through an autoregressive process. PVD Zhou et al. (2021) and 3DShape2VecSet Zhang et al. (2023a) are generative methods based on diffusion models. Point-MGE outperforms other methods on two of the evaluation metrics, indicating that our approach can generate 3D shapes that closely match the real distribution. Additionally, in Figure  2, we visualize the rendered results of the generated shapes.
Methods Airplane Faucet Table Chair
S-FID $\downarrow$ 3DILG 0.71 2.49 2.10 0.96
NeuralWavelet 0.38 1.84 1.12 1.14
3DShape2VecSet 0.62 1.47 1.19 0.76
Point-MGE 0.67 1.65 1.03 0.71
S-KID $\downarrow$ (×$\times 10^{3}$) 3DILG 0.81 3.76 3.84 1.21
NeuralWavelet 0.53 2.41 1.55 1.50
3DShape2VecSet 0.83 1.34 1.87 0.70
Point-MGE 0.79 1.31 1.14 0.58
Table 5: Category-conditional generation. We reported the S-FID and S-KID metrics for four categories.
@@
Extractor Generator Acc (%) R-FID $\downarrow$
6 6 91.2 29.07
8 6 91.9 27.85
12 8 92.4 21.63
12 12 92.1 20.94
Table 6: Effects of varying transformer depths. The accuracy (%) on the ScanObjectNN Uy et al. (2019) (OBJ_BG) and the R-FID for unconditional generation are reported.
@@
We evaluated the conditional generation capabilities of Point-MGE under two different settings. First, we retrained the model using the 54 categories from the ShapeNet Chang et al. (2015) dataset as conditional embeddings. As shown in Table 5, we evaluated the Surface-FID (S-FID) and Surface-KID (S-KID) for our method and the benchmark methods across four categories. The results indicate that our method achieved performance comparable to or even surpassing those of purely generative approaches. Figure 3 displays the rendered results of generated shapes for three categories (airplane, faucet, and table), with nine instances shown for each category. Next, we further evaluated Point-MGE against other SSL models in the masked generation task. As shown in Figure 4, these models typically train by reconstructing local point clouds, which primarily allows them to complete point cloud reconstruction tasks and depends on the provision of the central point positions. This approach, to some extent, leaks complete shape information to the decoder. In contrast, our method can reconstruct 3D shape meshes using the minimal amount of point cloud information and does not rely on masked center point information during the generation process.
Type Acc (%) R-FID $\downarrow$
Quantized features 92.4 21.63
Index-embedded tokens 87.2 20.34
Table 7: Effects of input type. The accuracy (%) and the R-FID are reported.
Sliding Masking Various Masking
$\beta=0.5$ $\beta=0.5$ $\beta=0.7$ $\mu=1.0$ $\mu=0.7$
$u=1$ $u=2$ $u=2$ $\sigma=0$ $\sigma=0.25$
Acc (%) 91.9 92.4 91.7 86.2 91.6
R-FID $\downarrow$ 24.25 21.63 22.71 26.72 22.43
Table 8: Effects of masking design.The accuracy (%) and the R-FID are reported.
NeRFs EMD CD L1 CD L2
Acc (%) 92.4 87.9 88.7 89.1
R-FID $\downarrow$ 21.63 N/A N/A N/A
Table 9: Effects of input type. The accuracy (%) and the R-FID are reported.
@@ 4.4 Ablation Studies
In this section, we discuss several factors that influence the representation learning and generative capabilities of the pre-trained model.
@@
As shown in Table 6, we used the standard single scale ViT architecture to build the extractor-generator structure and explored four different transformer depth combinations. The results indicate that increasing the model depth significantly enhances the performance of the pre-trained model. Additionally, consistent with trends observed in previous masked modeling studies He et al. (2022); Pang et al. (2022); Zhang et al. (2022b), an asymmetric extractor-generator structure is more beneficial for representation learning. This reliance on the extractor to obtain high-quality representations from a small number of visible tokens is due to the use of a relatively simple generator to complete the reconstruction task.
@@
In Table 7, we investigate the impact of two different input types on model performance. Specifically, after vector quantizing the initial features of the point cloud patches, we consider using either the quantized features or their indices in the codebook as inputs. Experimental results show that using index-embedded tokens benefits the generation task, but its generalization ability across datasets is limited, which is detrimental to fine-tuning on downstream tasks, leading to poorer representation learning capabilities.
@@
In determining the optimal masking strategy, as shown in Table 8, we compared our sliding masking strategy with the variable masking strategy used in MAGE Li et al. (2023) and analyzed the impact of different hyperparameters on performance. For the variable masking, $\mu$ and $\sigma$ represent the mean and standard deviation of the truncated Gaussian distribution, respectively. The results indicate that the variable masking ratio, which performs well on 2D images, does not provide an advantage in the point cloud modality, where data is relatively scarce. In contrast, our smooth masking strategy better facilitates the transition between representation learning and generative training.
Reconstruction Target.
In exploring the training of VQVAE, we compared different reconstruction targets. As shown in Table 9, this study employs NeRFs as the reconstruction target. Additionally, we experimented with directly using point clouds as the reconstruction target and tested methods using Earth Mover’s Distance (EMD) and Chamfer Distance (CD) as loss functions. The experimental results demonstrate that our approach excels in overcoming point cloud sampling bias, leading to superior representation learning capabilities. However, directly using point clouds as the reconstruction target, similar to previous methods, is limited to generating point clouds and cannot produce high-resolution 3D shapes. In the table, this generative capability is marked as Not Applicable (N/A).